# Improved Fine-Tuning of Large Multimodal Models for Hateful Meme Detection

### v1

@@ Abstract
Hateful memes have become a significant concern on the Internet, necessitating robust automated detection systems. While large multimodal models have shown strong generalization across various tasks, they exhibit poor generalization to hateful meme detection due to the dynamic nature of memes tied to emerging social trends and breaking news. Recent work further highlights the limitations of conventional supervised fine-tuning for large multimodal models in this context. To address these challenges, we propose Large Multimodal Model Retrieval-Guided Contrastive Learning (LMM-RGCL), a novel two-stage fine-tuning framework designed to improve both in-domain accuracy and cross-domain generalization. Experimental results on six widely used meme classification datasets demonstrate that LMM-RGCL achieves state-of-the-art performance, outperforming agent-based systems such as VPD-PALI-X-55B. Furthermore, our method effectively generalizes to out-of-domain memes under low-resource settings, surpassing models like GPT-4o.
@@ 1 Introduction
Previous studies focus on supervised settings, fine-tuning neural networks with in-domain training data Pramanick et al. (2021b); Kumar and Nandakumar (2022); Burbi et al. (2023); Lin et al. (2024). However, real-world challenges arise as memes continuously evolve with social trends and breaking news. This creates an out-of-domain generalization gap where these systems often fail to recognize new hate patterns Cao et al. (2024). Frequent retraining becomes impractical given daily content generation rates and associated annotation requirements. These challenges highlight the need for generalizable detection methods that perform well on in-domain examples while maintaining robustness in low-resource, out-of-domain scenarios Huang et al. (2024).
While large language models (LLMs) have demonstrated strong generalization across various tasks OpenAI (2023), recent studies indicate that vanilla Large Multimodal Models (LMMs) struggle with hateful meme detection, both in-domain and out-of-domain Mei et al. (2024); Huang et al. (2024); Hee et al. (2024); Cao et al. (2024). Notably, Mei et al. (2024) found that fine-tuned CLIP Radford et al. (2021) outperforms supervised fine-tuned (SFT) LMMs such as LLaVA-13B Liu et al. (2023b) and Flamingo-80B Alayrac et al. (2022), thus exposing shortcomings in standard SFT approaches for LMMs. Additionally, they showed that LMMs exhibit poor cross-domain generalization: for instance, a LLaVA model fine-tuned on the HatefulMemes dataset Kiela et al. (2021) (focusing on racist, sexist, and religion-based hate speech) fails to generalize to HarMeme Pramanick et al. (2021a), which targets COVID-related political memes. In addition, Huang et al. (2024) observed that hateful meme detection remains challenging for in-context learning framework in LMMs, suggesting that innovative approaches are needed to make better use of few-shot meme examples.
To address these challenges, we propose Large Multimodal Model Retrieval-Guided Contrastive Learning (LMM-RGCL), a two-stage fine-tuning framework designed to improve both in-domain and out-of-domain hateful meme detection. The first stage, Joint Multimodal Fine-tuning, rapidly adapts the LMM to the hateful meme detection task by integrating language modeling loss with cross-entropy loss for the classifier. The second stage, RGCL Fine-tuning, learns to discriminate between retrieved positive and negative examples, improving both classification and retrieval capabilities by aligning representations of semantically similar meme pairs. After LMM-RGCL fine-tuning, the learned embedding space enables retrieval-based KNN majority voting, improving classification performance on previously unseen examples. We show that this approach makes better use of few-shot meme examples than in-context learning. LMM-RGCL enables Qwen2VL-7B Wang et al. (2024) to achieve state-of-the-art performance on six popular meme classification datasets, surpassing larger agent-based systems such as VPD-PALI-X 55B Hu et al. (2024). Additionally, we show that LMM-RGCL fine-tuned models exhibit significantly stronger generalization to unseen meme datasets compared to standard SFT. Under low-resource settings, Qwen2VL-7B, using the retrieval-based KNN classifier, surpasses models such as GPT-4o OpenAI (2024a) and LOREHM with LLaVA-34B Liu et al. (2023a). Our contributions are:
@@
    We propose a novel two-stage fine-tuning framework, LMM-RGCL, which integrates joint multimodal fine-tuning with an RGCL fine-tuning phase. This approach optimizes both classification and retrieval capabilities for hateful meme detection.
@@
    We demonstrate that LMM-RGCL achieves state-of-the-art performance on six widely used meme classification datasets, outperforming larger agent-based systems that rely on standard supervised fine-tuning.
@@
    We show that LMM-RGCL significantly improves generalization to unseen meme datasets, offering a comprehensive solution for both in-domain and out-of-domain hateful meme detection. To our knowledge, this is the first system to attain state-of-the-art results in both scenarios.
@@ 2.1 Hateful Meme Detection
Most existing approaches to hateful meme detection use supervised learning. Early systems Zhu (2020); Muennighoff (2020) fine-tuned object detection–based vision-language models such as OSCAR Li et al. (2020), and UNITER Chen et al. (2020), which utilize Faster R-CNN Ren et al. (2015)-based object detectors Anderson et al. (2018); Zhang et al. (2021).
More recently, research has shifted toward CLIP Radford et al. (2021) for its end-to-end simplicity and stronger multimodal alignment. Numerous studies have fine-tuned models based on CLIP using different modality fusion mechanisms Pramanick et al. (2021b); Kumar and Nandakumar (2022); Shah et al. (2024). Other works incorporate caption models into the CLIP-based feature fusion network to further enhance performance Burbi et al. (2023); Cao et al. (2023); Ji et al. (2024). Additionally, contrastive learning techniques have been explored to address confounding factors in meme classification Lippe et al. (2020); Mei et al. (2024).
While LMMs such as Flamingo Alayrac et al. (2022) have shown promise in hateful meme detection via SFT, fine-tuning strategies for LMMs remain underexplored relative to CLIP-based approaches. In fact, Mei et al. (2024) demonstrated that fine-tuned CLIP models can outperform much larger LMMs, highlighting the need for specialized fine-tuning methods. In this work, we propose a novel fine-tuning approach for LMMs to improve their effectiveness in hateful meme detection.
@@ 2.2 Low resource hateful meme detection
Low-resource hateful meme detection has received relatively little attention, despite its growing importance in real-world deployments that require out-of-domain generalization. In this setting, an initially trained model is deployed to a new domain without gradient updates, relying only on demonstration examples for inference Huang et al. (2024). Hee et al. (2024) utilized text similarity–based few-shot examples to help LMMs generalize to unseen memes. Similarly, Hu et al. (2024) and Huang et al. (2024) explored agent-based LMM systems with few-shot learning for out-of-domain settings. However, Huang et al. (2024) observed that in-context learning is less effective for meme classification compared to other tasks, highlighting the need for more effective strategies to use demonstration examples.
Figure 1: Architecture of LMM-RGCL. We decompose the LMM into two components: the LMM Backbone and the LM Head (LMH). For each training example $i$, the last hidden state $\mathbf{h}_{i}$ is fed to the LMH to obtain the LM loss $\mathcal{L}_{i}^{LM}$. $\mathbf{h}_{i}$ is also fed to a trainable multilayer perceptron (MLP) to generate an embedding $\mathbf{g}_{i}$ for use as a retrieval query and as a feature for the Logistic Regression Classifier (LRC) to compute the cross entropy loss $\mathcal{L}_{i}^{LR}$. During training, pseudo-gold and hard negative examples are retrieved from the encoded meme database $\mathbf{G}$ for computing the contrastive loss $\mathcal{L}_{i}^{RGCLL}$. At inference, the same process retrieves the $K$ nearest neighbors for Retrieval-based KNN Classification (RKC), which predicts the label $\hat{y}_{t}^{RKC}$ for an inference example $t$.
In contrast, we use a retrieval-based majority voting scheme for classifying unseen memes and find that it makes more effective use of demonstration examples than conventional in-context learning.
3 LMM-RGCL Methodology
@@ 3.1 Preliminaries
3.2 LMM-RGCL Framework
Architecture As illustrated in Figure 1, LMM-RGCL integrates an LMM with two additional trainable components: a Multilayer Perceptron (MLP) that projects the LMM final hidden state $\mathbf{h}_{i}$ into an embedding $\mathbf{g}_{i}$ for use in classification and retrieval; and an LRC operating on $\mathbf{g}_{i}$ Figure 1 shows how the architecture supports multiple fine-tuning and inference modes.
Retrieval During training, FAISS-based Johnson et al. (2021) nearest neighbor search retrieves pseudo-gold positive Mei et al. (2024) and hard negative examples Schroff et al. (2015) for contrastive learning from the encoded meme database $\mathbf{G}$. At inference, FAISS is used to retrieve neighbors for the Retrieval-based KNN Classifier (RKC).
Inference modes Figure 1 shows three different classifiers: LMH, LRC, and RKC. For pre-trained and SFT LMMs, we report LMH results following prior work (Section 3.1). For LMM-RGCL models, we report the LRC results, unless otherwise specified. Section 4.5 presents a detailed comparison of the three inference modes.
3.3 Stage 1: Joint Multimodal Fine-tuning
@@
3.4 Stage 2: RGCL Fine-tuning
@@
$\mathcal{L}_{i}^{\textrm{Stage2}}=\mathcal{L}_{i}^{RGCLL}+\mathcal{L}_{i}^{LR},$ (5)
where $\mathcal{L}_{i}^{LR}$ is defined in Eq. 4, and $\mathcal{L}_{i}^{RGCLL}$ is the Retrieval-Guided Contrastive Learning Loss.
To compute $\mathcal{L}_{i}^{RGCLL}$, we retrieve pseudo-gold positive and hard negative examples from the training set. Specifically, for a given sample $i$ with embedding $\mathbf{g}_{i}$, we use FAISS Johnson et al. (2021) to perform the nearest neighbor search between $\mathbf{g}_{i}$ and every other target embedding $\mathbf{g}_{j}\in\mathbf{G}$ from the training set. The encoded meme database $\mathbf{G}$ is updated every 100 steps during fine-tuning.
Pseudo-gold positive examples are same-label examples that have high similarity scores with $\mathbf{g}_{i}$, while hard negative examples are opposite-label examples that have high similarity scores. We denote the embedding of the pseudo-gold positive example and hard negative example as $\mathbf{g}_{i}^{+}$ and $\mathbf{g}_{i}^{-}$, respectively. $\mathcal{L}_{i}^{RGCLL}$ is then computed as:
$\mathcal{L}_{i}^{RGCLL}$ $=L(\mathbf{g}_{i},\mathbf{g}_{i}^{+},\mathbf{g}_{i}^{-})$
@@
3.5 Retrieval Based KNN Classification
@@
Additionally, to enable RKC on pretrained or SFT LMMs that do not incorporate an MLP, we use the last hidden state $\mathbf{h}_{i}$ for the nearest neighbor search. The results are provided in Appendix D.
@@ 4 Experiments
We evaluate the performance of our systems on six popular meme classification datasets: HatefulMemes Kiela et al. (2021), HarMeme Pramanick et al. (2021a), MAMI Fersini et al. (2022), Harm-P Pramanick et al. (2021b), MultiOFF Suryawanshi et al. (2020) and PrideMM Shah et al. (2024). These datasets encompass varying definitions of harmful content (hateful, offensive, or targeted harassment) across different sociopolitical contexts. A detailed description and statistics are in Appendix A.
@@
16 w/ LMM-RGCL 91.1 82.1 93.2 88.1 90.4 79.9 91.6 91.1 71.1 64.8 78.1 78.4
[…]
Table 1: Comparison with baseline systems under supervised settings. For large multimodal models, we report the pre-trained models zero-shot and few-shot performance (using 4-shot evaluation), along with a comparison between SFT and LMM-RGCL. Additionally, for each LMM, we provide the  p-value from significance testing between SFT and LMM-RGCL. All  p-values are below 0.05. Best performance is highlighted in bold; second-best is underlined.
For HatefulMemes, HarMeme, and MAMI, we report the Area Under the Receiver Operating Characteristic Curve (AUC) and Accuracy (Acc) in line with previous studies Kumar and Nandakumar (2022); Cao et al. (2023); Mei et al. (2024); Cao et al. (2024). For Harm-P, MultiOFF, and PrideMM, we report Accuracy and F1 score consistent with the literature Pramanick et al. (2021b); Mei et al. (2024); Shah et al. (2024); Lin et al. (2024). Implementation details, including hyperparameters, and statistical significance test procedures are detailed in Appendix B.
4.1 Comparing LMM-RGCL to Baseline Systems under Supervised Settings
Table 1 presents the performance of baseline systems under supervised fine-tuning settings. We compare LMM-RGCL against a range of strong baselines: the best prior models for each dataset [^1] From a recent paper Nguyen and Ng (2024); some datasets have been updated with the new best results.; supervised fine-tuned CLIP-based classifiers; and Large Multimodal Models (LMMs). All models are fine-tuned and evaluated for each dataset separately.
CLIP-based Classifiers We compare the performance of fine-tuned CLIP Radford et al. (2021) model with three other fine-tuning methods for CLIP-based systems: MOMENTA Pramanick et al. (2021b), HateCLIPper Kumar and Nandakumar (2022) and RGCL Mei et al. (2024).
Large Multimodal Models We experiment with three LMMs from two model families: LLaVA-1.5-7B Liu et al. (2023a), Qwen2VL-2B and Qwen2VL-7B Wang et al. (2024). We report the performance of these LMMs in the following settings: pre-trained models with zero-shot and few-shot prompts using the LMH; SFT LMMs using the LMH; and classification using LRC under the LMM-RGCL fine-tuning framework.
Best Prior Models Visual Program Distillation (VPD) Hu et al. (2024) and ExplainHM Lin et al. (2024) are LLM agent-based systems. The remaining state-of-the-art models, including ISSUES Burbi et al. (2023), Pro-Cap Cao et al. (2023), RGCL Mei et al. (2024) and MemeCLIP Shah et al. (2024), are based on fine-tuning CLIP-based vision and language models. Detailed descriptions of these methods are provided in Appendix C.

### v2

# Robust Adaptation of Large Multimodal Models for Retrieval Augmented Hateful Meme Detection

@@ Abstract
Hateful memes have become a significant concern on the Internet, necessitating robust automated detection systems. While LMMs have shown promise in hateful meme detection, they face notable challenges like sub-optimal performance and limited out-of-domain generalization capabilities. Recent studies further reveal the limitations of both SFT and in-context learning when applied to LMMs in this setting. To address these issues, we propose a robust adaptation framework for hateful meme detection that enhances in-domain accuracy and cross-domain generalization while preserving the general vision-language capabilities of LMMs. Experiments on six meme classification datasets show that our approach achieves state-of-the-art performance, outperforming larger agentic systems. Moreover, our method generates higher-quality rationales for explaining hateful content compared to standard SFT, enhancing model interpretability.
Robust Adaptation of Large Multimodal Models for Retrieval Augmented Hateful Meme Detection
Jingbiao Mei, Jinghong Chen, Guangyu Yang, Weizhe Lin, Bill Byrne Department of Engineering University of Cambridge Cambridge, United Kingdom, CB2 1PZ {jm2245, jc2124, gy266, wl356, wjb31}@cam.ac.uk
@@ 1 Introduction
Large Multimodal Models (LMMs) have emerged as a promising solution for this complex task Hee et al. (2024b); Lin et al. (2025). Their strong capabilities across a range of general vision-language tasks provide a solid foundation for understanding the intricate interplay between text and image in memes Zhu et al. (2023); Liu et al. (2023b). Furthermore, the generative nature of LMMs offers interpretability, allowing models to provide rationales for their detection decisions. Ideally, LMMs also bring improved generalizability, enabling them to adapt to the rapidly evolving landscape of online memes, making them well-suited for deployment in real-world content moderation systems. Despite this potential, current LMMs face the following challenges when applied to hateful meme detection.
@@
    Sub-optimal performance. LMMs struggle to learn the interplay of visual and textual cues inherent in hateful memes through standard supervised fine-tuning (SFT), as reported by Mei et al. (2024). We also report that SFT LMMs produce lower-quality rationales when explaining hateful content, possibly caused by overfitting and the scarcity of the training data.
@@
    Limited out-of-domain generalization. Memes constantly evolve with social trends and events, posing a generalization challenge Cao et al. (2024); Mei et al. (2024). While in-context learning with retrieved examples from a dynamic meme database is a potential approach to generalize to unseen data for LMMs, Huang et al. (2024) found that this approach remains ineffective, highlighting the need for more effective methods to use few-shot meme examples.
@@
    Degradation of general vision-language abilities can arise from fine-tuning for meme classification. We observe that applying SFT for meme classification leads to overfitting, which degrades performance on general multimodal benchmarks like MMMU Yue et al. (2023). This undermines the rationale for choosing LMMs over single-purpose specialized models such as CLIP.
14%25%61%SFTTieRA-HMD
Figure 1: Comparison of rationales generated by SFT and RA-HMD Qwen2VL-7B models on the HatefulMemes dataset. The bar chart shows the winning rate of rationale quality based on pairwise comparisons between the two models. A more detailed analysis is provided in Appendix K.
To address these challenges, we propose RA-HMD (Retrieval-Augmented Hateful Meme Detection), a framework that incorporates architectural enhancements and a two-stage fine-tuning strategy to adapt LMMs for hateful meme detection without degradation in general vision-language ability.
We address the three challenges of applying LMMs to hateful meme classification through the following contributions:
  1. 1.
    We propose RA-HMD, a fine-tuning framework for adapting LMMs for hateful meme classification, achieving new state-of-the-art results on six widely used meme classification datasets. In addition, RA-HMD generates higher quality rationales compared to SFT models, thereby enhancing the interpretability of LMM predictions, as shown in Figure 1.
  2. 2.
    RA-HMD demonstrates more robust out-of-domain generalization compared to SFT models. Notably, when RA-HMD is combined with a retrieval-augmented KNN classifier, it demonstrates state-of-the-art performance for out-of-domain meme classification. This setup leverages few-shot meme examples more effectively than in-context learning, addressing the challenge of adapting to rapidly evolving memes without retraining.
  3. 3.
    RA-HMD expands LMMs ability to perform hateful meme classification and explaining hateful memes without compromising performance on other vision–language tasks, as shown by results in Section 4.4.
@@ 2.1 Hateful Meme Detection
Most existing approaches to hateful meme detection rely on supervised learning, with the majority of research leveraging CLIP Radford et al. (2021). Numerous studies have fine-tuned models based on CLIP using different modality fusion mechanisms Pramanick et al. (2021b); Kumar and Nandakumar (2022); Shah et al. (2024). Other works incorporate caption models into the CLIP-based feature fusion network to further enhance performance Burbi et al. (2023); Cao et al. (2023); Ji et al. (2024). Additionally, contrastive learning techniques have been explored to address confounding factors in meme classification Lippe et al. (2020); Mei et al. (2024).
With the emergence of LMMs, recent research has shifted toward using LMMs as generalist models, in contrast to the specialist nature of CLIP based models Laurençon et al. (2023); Hu et al. (2024). Moreover, decoder-based LMMs offer an additional advantage: they can generate textual rationales to explain why a meme may be hateful Lin et al. (2024); Hee et al. (2024b).
While LMMs such as Flamingo Alayrac et al. (2022) have shown promise in hateful meme detection via SFT, fine-tuning strategies for LMMs remain underexplored. In fact, Mei et al. (2024) demonstrated that fine-tuned CLIP models can outperform much larger LMMs, highlighting the need for specialized methods. In this work, we address this gap by proposing LMM architecture refinement alongside a novel fine-tuning approach for LMMs that enhances their performance on hateful meme detection while preserving their general vision–language capabilities.
@@ 2.2 Low resource hateful meme detection
Low-resource hateful meme detection is critical for real-world applications that demand out-of-domain generalization. In this setting, an initially trained model is deployed to a new domain without gradient updates, relying only on demonstration examples for inference Huang et al. (2024). Hee et al. (2024a) utilized retrieved few-shot examples to help LMMs generalize to unseen memes. Hu et al. (2024) and Huang et al. (2024) explored agent-based LMM systems with few-shot learning for out-of-domain settings. However, Huang et al. (2024) observed that in-context learning is less effective for meme classification compared to other tasks, highlighting the need for more effective strategies to use demonstration examples.
Figure 2: Architecture of RA-HMD. We decompose the LMM into two components: the LMM Backbone and the LM Head (LMH). For each training example $i$, the last hidden state $\mathbf{h}_{i}$ is fed to the LMH to obtain the LM loss $\mathcal{L}_{i}^{LM}$. $\mathbf{h}_{i}$ is also fed to a trainable multilayer perceptron (MLP) to generate an embedding $\mathbf{g}_{i}$ for use as a retrieval query and as a feature for the Logistic Regression Classifier (LRC) to compute the cross entropy loss $\mathcal{L}_{i}^{LR}$. During training, contrastive learning examples are retrieved from the encoded meme database $\mathbf{G}$ for computing the contrastive loss $\mathcal{L}_{i}^{CL}$. At inference, the same process retrieves the $K$ nearest neighbors for Retrieval-augmented KNN Classification (RKC), which predicts the label $\hat{y}_{t}^{RKC}$ for an inference example $t$.
In this work, we show that RA-HMD improves the in-context learning capabilities of LMMs. Furthermore, when combined with a retrieval-augmented KNN classifier, RA-HMD enables more effective use of demonstration examples than conventional in-context learning.
3 RA-HMD Methodology
@@ 3.1 Preliminaries
3.2 RA-HMD Framework
Architecture enhancement Leveraging representations from large multimodal models (LMMs) for hateful meme classification is non-trivial, particularly when attempting to use LMM embeddings for classification while preserving the model’s original language generation capabilities. Prior work has explored various strategies for adapting these representations to classification and retrieval tasks. In our study, we similarly experimented with multiple adaptation methods. Appendix D provides a comprehensive summary of these efforts, including failure cases of previous approaches and key insights that ultimately guided the design of RA-HMD.
As illustrated in Figure 2, RA-HMD integrates an LMM with two additional trainable components: a Multilayer Perceptron (MLP) that projects the LMM final hidden state $\mathbf{h}_{i}$ into an embedding $\mathbf{g}_{i}$ for use in classification and retrieval; and an LRC operating on $\mathbf{g}_{i}$ Figure 2 shows how the architecture supports multiple fine-tuning and inference modes.
Retrieval During stage-2 training, FAISS-based Johnson et al. (2021) nearest neighbor search retrieves contrastive learning examples from the encoded meme database $\mathbf{G}$. At inference, FAISS is used to retrieve neighbors for the Retrieval-augmented KNN Classifier (RKC).
Inference modes Figure 2 shows three different classifiers: LMH, LRC, and RKC. For pre-trained and SFT LMMs, we generate classification decisions using the LMH as described in Section 3.1. For RA-HMD models, we obtain meme classification decisions via the LRC, unless otherwise specified. Section 4.7 presents a detailed comparison of the three inference modes.
3.3 Stage 1: Logistic Regression Augmented Supervised Fine-tuning
@@
3.4 Stage 2: LMM Contrastive Fine-tuning
@@
$\mathcal{L}_{i}^{\textrm{Stage2}}=\mathcal{L}_{i}^{CL}+\mathcal{L}_{i}^{LR},$ (5)
where $\mathcal{L}_{i}^{LR}$ is defined in Eq. 4, and $\mathcal{L}_{i}^{CL}$ is the Contrastive Learning Loss.
To compute $\mathcal{L}_{i}^{CL}$, we retrieve pseudo-gold positive examples similar to RGCL Mei et al. (2024) and hard negative examples Schroff et al. (2015) from the training set. Specifically, for a given sample $i$ with embedding $\mathbf{g}_{i}$, we use FAISS Johnson et al. (2021) to perform the nearest neighbor search between $\mathbf{g}_{i}$ and every other target embedding $\mathbf{g}_{j}\in\mathbf{G}$ from the training set.
Pseudo-gold positive examples are same-label examples that have high similarity scores with $\mathbf{g}_{i}$, while hard negative examples are opposite-label examples that have high similarity scores. We denote the embedding of the pseudo-gold positive example and hard negative example as $\mathbf{g}_{i}^{+}$ and $\mathbf{g}_{i}^{-}$, respectively. $\mathcal{L}_{i}^{CL}$ is then computed as:
$\mathcal{L}_{i}^{CL}$ $=L(\mathbf{g}_{i},\mathbf{g}_{i}^{+},\mathbf{g}_{i}^{-})$
@@
3.5 Retrieval Augmented KNN Classification
@@
Additionally, to enable RKC on pretrained or SFT LMMs that do not incorporate an MLP, we use the last hidden state $\mathbf{h}_{i}$ for the nearest neighbor search. The results are provided in Appendix F.
@@ 4 Experiments
We evaluate on six meme classification datasets: HatefulMemes Kiela et al. (2021), HarMeme Pramanick et al. (2021a), MAMI Fersini et al. (2022), Harm-P Pramanick et al. (2021b), MultiOFF Suryawanshi et al. (2020) and PrideMM Shah et al. (2024). A detailed description and statistics are in Appendix A. Implementation details are described in Appendix B.
@@
16 w/ RA-HMD 91.1 82.1 93.2 88.1 90.4 79.9 91.6 91.1 71.1 64.8 78.1 78.4
Table 1: Comparison with baseline systems under supervised settings. For large multimodal models, we report the pre-trained models zero-shot and few-shot performance (using 4-shot evaluation), along with a comparison between SFT and RA-HMD. Best performance is highlighted in bold; second-best is underlined.
4.1 Comparing RA-HMD to Baseline Systems under Supervised Settings
Table 1 presents the performance of baseline systems under supervised fine-tuning settings. We compare RA-HMD against a range of strong baselines: the best prior models for each dataset [^1] From a recent paper Nguyen and Ng (2024); some datasets have been updated with the new best results.; supervised fine-tuned CLIP-based classifiers; and Large Multimodal Models (LMMs). All models are fine-tuned and evaluated for each dataset separately.
CLIP-based Classifiers We compare the performance of fine-tuned CLIP Radford et al. (2021) model with two other fine-tuning methods for CLIP-based systems: HateCLIPper Kumar and Nandakumar (2022) and RGCL Mei et al. (2024).
Large Multimodal Models We experiment with three LMMs from two model families: LLaVA-1.5-7B Liu et al. (2023a), Qwen2VL-2B and Qwen2VL-7B Wang et al. (2024b). We report the performance of these LMMs in the following settings: pre-trained models with zero-shot and few-shot prompts using the LMH; SFT LMMs using the LMH; and classification using LRC under the RA-HMD fine-tuning framework. We further include the results with GPT-4o OpenAI (2024a) with optimized prompting for each dataset for reference. For GPT-4o, the token likelihood is not accessible to compute the AUC score.
Best Prior Models Visual Program Distillation (VPD) Hu et al. (2024) and ExplainHM Lin et al. (2024) are LLM agent-based systems. The remaining state-of-the-art models, including ISSUES Burbi et al. (2023), Pro-Cap Cao et al. (2023), RGCL Mei et al. (2024) and MemeCLIP Shah et al. (2024), are based on fine-tuning CLIP-based vision and language models. Detailed descriptions of these methods are provided in Appendix M.